Calculate pFe2+pFe2+ at each of the points in the titration of 25.00 mL of 0.020580.02058 M Fe2+Fe2+ by 0.036580.03658 M EDTAEDTA at a pH of 7.007.00. The values for log?flog⁡Kf and ?Y4−αY4− can be found in the chempendix.

log Kf=14.30

\alphay4-= 3.8x10^-4

a. 11.50 mL . pFe2+=

b. the equivalence point, ?e

c.18.50 mL . pFe2+=
Answer #1


14.30 Kg = 100*xxy = 1014-30 * 3.87164 = 7.58 x100 Ve X 0.03658 = 25X0.02058 9) re=14.065 [Fe2+] =(14.065- 11:50 1940.02058*